A perimeter of a rectangle is 8 meters and its area is 3.84 m. What are the dimensions?
Answer by edjones(8007)

2L+2w=P
2L+2w=8
L+w=4
L=4-w
.
Lw=A
Lw=3.84
w(4-w)=3.84
4w-w^2=3.84
w^2-4w=-3.84 multiply each side by -1.
w^2-4w+4=4-3.84 completing the square.
(w-2)^2=.16
w-2=+-.4
w=1.6m, L=2.4m
.
Ed